OpenStudy (anonymous):

a line is prepindicular to the line y=2x+3 and has the same x-intercept as x+3y+10=0. find the equation for this line in the y=mx+b form!

OpenStudy (anonymous):

As for your question: To find the x-intercept of the second equation. Let y=0. x+10=0 x=-10. Then perpendicular lines have an inverse reciprocal slope. So 2 becomes -1/2. So: y=(-1/2)x-5 (because if you let y=0 you can see the x-intercept is -10)